### v1

# Dynamic Decoupling in Multidimensional Screening

Abstract
I study multidimensional sequential screening. A monopolist contracts with an agent endowed with private information about the distribution of their eventual valuations of different goods; a contract is written and the agent reports their initial private information before drawing and reporting their valuations. In these settings, the monopolist frontloads surplus extraction: Any information rents given to the agent to elicit their post-contractual valuations can be extracted in expectation before valuations are drawn. This significantly simplifies the multidimensional screening problem. If the agent’s valuations satisfy invariant dependencies (valuations can be dependent across dimensions, but how valuations are coupled cannot vary in their initial private information), the optimal mechanism coincides with independently offering the optimal sequential screening mechanism for each good, regardless of the dependency structure.
Keywords: Dynamic Contracting, Dynamic Irrelevance, Multidimensional Screening.
JEL Codes: D47.
1 Introduction
Multidimensional screening and dynamic contracting problems are generally difficult to solve: A multi-good monopolist’s revenue-maximizing mechanism often features nonlinear bundling and uncountably large menus. In this light, why does Amazon not bundle products? The answer is Amazon Prime: Extracting surplus intertemporally is more effective than at the moment of sale and simplifies the principal’s problem.
I study a model of dynamic multidimensional screening where a monopolist contracts with an agent who arrives with (one-dimensional) private information about their eventual distribution over (multidimensional) valuations; the model is a straightforward extension of CourtyLi2000 to multiple goods. Endowing the principal with the ability to contract intertemporally smooths out interdimensional incentive constraints, leading to separately selling different goods being optimal in a large variety of settings. I formalize a new force in dynamic adverse selection problems where the agent’s information arrives over time: The principal prefers to frontload surplus extraction, analogous to how transfers are backloaded in dynamic moral hazard problems. There are two forces, one governing feasibility and one governing optimality, for why surplus extraction is frontloaded. First is feasibility, coming from a pushforward force [^2] See, for example, the pullback lemma of luo2025payingpersuading.: Before the agent’s valuations are realized, the principal can extract the expected information rents the agent receives after their valuations are realized. Second is optimality. The principal needs to provide less information rents to the agent when the agent simply has less private information: Joint deviations have less bite as current period deviations can only be coordinated with future deviations in expectation. As such, the only information rents that need to actually be paid to the agent are those needed to elicit the agent’s pre-contractual information, in line with the dynamic irrelevance results of EsoSzentes2017. Post-contractual information rents can be provided “for free.”
Instead of providing the agent with their full static information rents, pulling surplus extraction forward distorts the agent’s information rents by (1) the inverse hazard ratio for their pre-contractual information and (2) the score of a likelihood ratio test for their valuations being drawn from the distribution corresponding to their report. This distortion not only works in favor of the principal, but additionally significantly simplifies the screening problem when the joint distribution of pre-contractual information and valuations satisfies invariant dependencies, generalizing conditional independence. Different dimensions of the agent’s valuations can be dependent, but the structure of these dependencies cannot vary in the agent’s pre-contractual private information. When this is the case, a likelihood ratio test is separable in each dimension and selling each good separately is optimal. The optimal contract has the same structure as CourtyLi2000: The agent is initially offered a menu of upfront payments and option payments for each good to be exercised when their valuation is drawn in the future. This mirrors real-world memberships at large retailers, such as Amazon Prime.
Finally, I compare my model and notion of dynamic irrelevance to other models of dynamic contracting. First, in an alternative setting where the agent once again starts with the same pre-contractual private information but valuations are revealed and goods are sold in sequence instead of simultaneously (as in segal2014), if valuations are independent then the monopolist’s optimal mechanism and profit are the same as the baseline case. Second, in the relaxed problem where an orthogonalized signal of the agent’s valuations were publicly observed (as in EsoSzentes2017) the monopolist’s optimal mechanism and profit are once again unchanged.
2 Model
The model is a straightforward extension of CourtyLi2000 to a multidimensional setting. A monopolist contracts with a single buyer. The monopolist sells $n$ goods. Before contracting, the buyer observes $\gamma\in[\underline{\gamma},\overline{\gamma}]=\Gamma$ drawn according to $G(\gamma)$. After contracting, the buyer draws $\theta\in\Theta=\prod_{j}[\underline{\theta}^{j},\overline{\theta}^{j}]\subseteq\mathbb{R}^{n}$ according to $F(\theta|\gamma)$. I assume $F,G$ have densities $f,g$ and $f$ is continuously differentiable with respect to $\gamma$ and $\theta$. Notationally, let superscripts denote marginals and subscripts denote partial derivatives, so $F^{j}_{\gamma}$ would denote the partial derivative of the marginal of $F$ in the $j$th dimension with respect to $\gamma$. For regularity, suppose the usual one-dimensional regularity condition holds marginal-by-marginal: for each marginal $j$, we have that $F^{j}_{\gamma}(\theta^{j}|\gamma)\leq 0$ and “virtual value”
$\theta^{j}+\frac{F_{\gamma}^{j}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}$
is increasing in $\gamma$ and single-crossing in $\theta^{j}$.
The monopolist chooses $q:\Gamma\times\Theta\to[0,1]^{n},t_{1}:\Gamma\to\mathbb{R},t_{2}:\Gamma\times\Theta\to\mathbb{R}$. The buyer chooses whether to participate in the mechanism and what $\gamma^{\prime}$ to report if they do participate after observing $\gamma$ (but before observing $\theta$). Then, the buyer observes $\theta$ and decides what $\theta^{\prime}$ to report. The monopolist’s payoff is $t_{1}(\gamma^{\prime})+t_{2}(\gamma^{\prime},\theta^{\prime})$. The buyer’s payoff from receiving a (probabilistic) vector of goods $q$ while paying $t_{1},t_{2}$ is
$q\cdot\theta-t_{1}-t_{2}$
where $\cdot$ is the dot product, so the buyer’s value is additive across goods, there is no discounting across time, and $\gamma$ is payoff-irrelevant (except for determining the distribution of valuations). By the Revelation Principle, the monopolist’s problem is:
$\max_{t_{1},t_{2},q}$ $\>\mathbb{E}_{\gamma,\theta}\left[t_{1}(\gamma)+t_{2}(\gamma,\theta)\right]$
s.t. $\>\gamma\in\operatorname*{\arg\!\max}_{\gamma^{\prime}}\left\{\mathbb{E}_{\theta|\gamma}\left[\max_{\theta^{\prime}}q(\gamma^{\prime},\theta^{\prime})\cdot\theta-t_{1}(\gamma^{\prime})-t_{2}(\gamma^{\prime},\theta^{\prime})\right]\right\}\text{ for all }\gamma\in\Gamma;$ $IC1$
$\>\theta\in\operatorname*{\arg\!\max}_{\theta^{\prime}}q(\gamma,\theta^{\prime})\cdot\theta-t_{2}(\gamma,\theta^{\prime})\text{ for all }\theta\in\Theta;$ $IC2$
$\>\mathbb{E}_{\theta|\gamma}\left[q(\gamma,\theta)\cdot\theta-t_{1}(\gamma)-t_{2}(\gamma,\theta)\right]\geq 0\text{ for all }\gamma\in\Gamma.$ $IR$
The constraints are as follows. First, $IC1$ pins down the buyer having no incentive to misreport $\gamma$ for any possible continuation misreport of $\theta$ after their new information arrives. However, their value of $\gamma$ still pins down the distribution of $\theta$ they face tomorrow. Next, $IC2$. After $\gamma$ is (truthfully) revealed, the buyer has no incentive to misreport $\theta$. Finally, $IR$ is an interim participation constraint: The buyer’s choice of whether or not to participate in the contract can depend on their observed value of $\gamma$.
3 Dynamic (Ir)relevance
We first present a rewriting of the problem that simplifies the monopolist’s problem and illustrates the role of frontloading surplus extraction.
Lemma 1.
The monopolist’s problem can be written as
$\max_{u(\gamma,\theta),t_{1}(\gamma)}\text{ }$ $\mathbb{E}_{\gamma,\theta}\left[\theta\cdot\nabla_{\theta}u(\gamma,\theta)-u(\gamma,\theta)\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}\right]$
st. $\gamma\in\operatorname*{\arg\!\max}\mathbb{E}_{\theta|\gamma}\left[u(\gamma^{\prime},\theta)\right]-t_{1}(\gamma^{\prime})$ $IC1$
$u(\gamma,\theta)\text{ is convex in }\theta;$ $IC2$
$u(\gamma,\theta)\text{ is 1-Lipschitz in }\theta;$ $F$
$\mathbb{E}_{\theta|\gamma}\left[u(\gamma,\theta)\right]-t_{1}(\gamma)\geq 0\text{ for all }\gamma$ $IR$
where $u(\gamma,\theta)=\theta\cdot q(\gamma,\theta)-t_{2}(\gamma,\theta)$ is the agent’s utility net of $t_{1}$ and $\nabla_{\theta}u(\gamma,\theta)$ is the gradient of $u(\gamma,\theta)$ in $\theta$ at some fixed $\gamma$.
$IC1$ continues to pin down incentive compatibility for truthful reporting of $\gamma$ while $IC2$ captures incentive compatibility for truthful reporting of $\theta$. Next, $F$ is feasibility: The gradient of the agent’s post-contractual utility function $u$ is equal to the allocation rule which must lie between zero and one. Finally, $IR$ pins down interim individual rationality to accept the contract. Total surplus generated by the monopolist is $\theta\cdot\nabla_{\theta}u(\gamma,\theta)$ while a share $u(\gamma,\theta)\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}$ is paid to the agent as information rents. The Lemma establishes that the interim utility function $u$ pins down both period one and period two transfers and specifies how dynamics distorts designer profits.
Compared to the usual multi-good monopolist problem where the agent receives their full information rents $u$, allowing the principal to frontload surplus extraction leads to period two information rents given to the agent being extracted by period one transfers. [^3] In the proof of the Lemma, we explicitly work out how the cancellation happens. As such, the only information rents the principal needs to pay the agent are those to truthfully elicit their pre-contractual private information. These information rents are distorted by two factors. First, $\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}=\frac{\partial}{\partial\gamma}\ln(f(\theta|\gamma))=s(\gamma,\theta)$ is the score function measuring how much the log-likelihood of $\theta$ changes as $\gamma$ changes. This captures how likely $\theta$ came from one value of $\gamma$ versus another and thus governs how post-contractual utilities are passed through into information rents. The other $\frac{1-G(\gamma)}{g(\gamma)}$ term is the standard hazard rate; as usual, providing information rents to lower types once again leads to higher information rents to higher types.
Applying the divergence theorem as in opttransport allows us to express $\nabla_{\theta}u(\gamma,\theta)$ in terms of $u(\gamma,\theta)$. However, this significantly complicates the expression and does not offer much additional insight; in fact, it may be simpler to express the problem in terms of $\nabla_{\theta}u(\gamma,\theta)$ instead. There are two further complications preventing us from using an optimal transport or duality (see kleiner2019) approach to analyze the problem. First, as segal2014 notes, the consumer’s expected interim utility function may be discontinuous in $\gamma$. Unlike continuity in $\theta$ being guaranteed by feasibility of the allocation rule, the principal may choose to give information rents to consumers with different values of $\gamma$ at different values of $\theta$. While is is possible to use an orthogonalization approach to transform the problem, making $\mathbb{E}_{\theta|\gamma}\left[u(\gamma,\theta)\right]$ continuous in $\gamma$, it is unclear if $u(\gamma,\theta)$ itself can be made continuous in $\gamma$; or if doing so would be an useful exercise. Second, the usual single-crossing conditions that lead to monotonicity being a necessary and sufficient condition for implementability do not hold. We can view the problem as the principal designing a utility function $w$ where an agent with private information $\gamma$ values $w$ at
$\hat{U}(\gamma,w(\theta))=\int_{\Theta}w(\theta)dF(\theta|\gamma).$
Then, the usual characterizations of single crossing or monotonicity in $w$ cannot be applied as $w$ is a high-dimensional object. However, cyclic monotonicity as in rochet1987 is still a necessary and sufficient condition for implementability, but is difficult to operationalize.
When does Lemma 1 help? Consider the case when different dimensions of $\theta$ are independent (the following intuition generalizes): Then, multidimensional sequential screening problems have a simple solution. In this case, additivity of scores $s(\gamma,\theta)$ leads to independently selling each good being optimal: Each dimension of $\theta$ provides information about $\gamma$ separately. The joint forces of intertemporal screening and independence leads to the solution being separable: Intertemporal screening enables the principal to push surplus extraction to the first period while additivity of scores enables information rents to be computed good-by-good.
Formally, let $\mathbf{V}(\theta|\gamma)$ be a vector-valued function satisfying
$\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))=f_{\gamma}(\theta|\gamma).$ (1)
Then, applying the Divergence Theorem (when boundary conditions goes to zero) gives
$\int_{\Theta}u(\gamma,\theta)\left(\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))\right)d\theta=-\int_{\Theta}\nabla_{\theta}u(\gamma,\theta)\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))d\theta$
so the objective can be written as
$\mathbb{E}_{\gamma,\theta}\left[\nabla_{\theta}u(\gamma,\theta)\cdot\left(\theta+\mathbf{V}(\theta|\gamma)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$ (2)
In the independent case, $\mathbf{V}(\theta|\gamma)$ has a simple form: The $j$th component of $\mathbf{V}$ is
$\mathbf{V}^{j}(\theta|\gamma)=\frac{F^{j}_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}.$ (3)
This solution to $\mathbf{V}$ works in more general settings than perfect conditional independence. By Sklar’s Theorem, there exists an unique distribution $C:[0,1]^{n}\times\Gamma\to[0,1]$ such that
$F(\theta|\gamma)=C(F^{1}(\theta^{1}|\gamma),...,F^{n}(\theta^{n}|\gamma)|\gamma)$
and letting $c$ be the density of $C$,
$f(\theta|\gamma)=c(F^{1}(\theta^{1}|\gamma),...,F^{n}(\theta^{n}|\gamma)|\gamma)\prod_{j}f^{j}(\theta^{j}|\gamma).$
Definition 1 (Invariant Dependencies).
Distribution $F(\theta|\gamma)$ has invariant dependencies if $c(\cdot|\gamma)$ is constant in $\gamma$. If this is the case, we write $C(\cdot|\gamma)=C(\cdot),c(\cdot|\gamma)=c(\cdot)$.
If the dimensions of $F$ are independent conditional on $\gamma$, then $C$ is uniform over $0,1]^{n}$ for each $\gamma$ so $F$ has invariant dependencies. As such, this condition is a generalization of conditional independence; any choice of marginals $F^{1}(\theta^{1}|\gamma),...,F^{n}(\theta^{1}|\gamma)$ along with a $\gamma$-invariant non-uniform choice of $C$ over $0,1]^{n}$ produces a joint distribution $F$ that is not conditionally independent but satisfies invariant dependencies. In terms of percentiles, the marginals $F^{j}$ governs how percentiles in each dimension are mapped into valuations in each dimension while $C$ governs the dependency structure between percentiles. As long as the dependencies between percentiles is invariant as $\gamma$ changes, any $\gamma$-dependent mapping from percentiles to valuations for each dimension satisfies invariant dependencies.
If $F$ has invariant dependencies then the $\mathbf{V}$ identified in (3) satisfies condition (1) and only the marginals matter. Given regularity along each marginal, pointwise maximization good-by-good solves the problem and produces the same optimal mechanism regardless of the dependency structure between goods as long as that dependency is invariant in $\gamma$.
Theorem 1 (Invariant Dependencies Implies Separability).
If $F$ has invariant dependencies, then the optimal joint mechanism is to offer the optimal sequential screening mechanism for each good separately.
In the terminology of segal2014, $V$ is a vector of impulse responses of $\theta$ to $\gamma$. Recalling that $\nabla_{\theta}u(\gamma,\theta)=q(\gamma,\theta)$ gives the problem to be
$\mathbb{E}_{\gamma,\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\phi^{j}(\gamma,\theta^{j})\right]\text{ for }\phi^{j}(\gamma,\theta^{j})=\theta^{j}+\frac{F_{\gamma}^{j}(\theta^{j}|\gamma)}{f(\theta^{j}|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}.$
Here, $\phi^{j}$ coincides with the virtual value for good $j$ in a one-good sequential screening setting. When regularity holds, pointwise maximization produces a feasible solution; this pointwise maximization can be done good-by-good, leading to Theorem 1. Good-by-good the optimal mechanism consists of an upfront payment and an option value to buy; higher $\gamma$ consumers pay more upfront but have lower strike prices. Aggregating the optimal mechanism good-by-good leads to the optimal mechanism in a multidimensional screening setting to be one where the agent once again faces a menu of upfront payments (equal to the sum of upfront payments coming from the optimal mechanism for each good) and a separate strike price for each good.
3.1 Additional Periods
If the optimal mechanism is to simply sell goods independently, do dynamics have bite? The answer is no. Consider an alternative environment where instead of buying $n$ products simultaneously, there were $n+1$ periods. In the first period, the agent draws $\gamma$ and the monopolists writes a contract. In each subsequent period $i$ the buyer’s valuation $\theta^{i}$ is drawn according to $F^{i}(\theta^{i}|\gamma)$ and good $i$ is sold via some mechanism. Call this the sequential case. In general, profits from the original problem of selling simultaneously are upper bounded by profits from the alternative case of selling sequentially: When selling simultaneously, the buyer draws all of their valuations at once and is able to better coordinate joint deviations across goods. As such, any incentive compatible mechanism in the original problem is also incentive compatible in the sequential case, implying that the value of the sequential case is lower-bounded by the value of the original problem. However, this bound is tight whenever values are independent: Endowing the designer with additional power to screen intertemporally does not increase profits. Since all information rents are pulled forward into period one anyways, additional periods beyond period two are inconsequential. Note that this result is not a direct consequence of EsoSzentes2017: Dynamic irrelevance in their setting only pertains to not needing to pay information rents to elicit future information so the same allocation rule leads to the same profit. I contribute by showing that even when expanding the space of implementable allocation rules, the principle does not do better when screening additionally through time. As such, the role of $\gamma$ (learning about learning) is qualitatively different than $\theta$ (learning about valuations).
Proposition 1 (Irrelevance of Additional Periods).
If values are independent, then the optimal allocation rule and profit from selling simultaneously coincides with the optimal allocation rule and profit from selling sequentially.
If there is dependence and different dimensions are drawn sequentially, however, different components of $\theta$ leads to both learning about learning and learning about valuations. The proof of Proposition 1 compares the solution to a multidimensional sequential screening problem to the optimal mechanism identified in segal2014; absent independence, the optimal mechanism in the sequential case is difficult to derive.
3.2 Observed Orthogonalization
The standard approach to solving dynamic contracting problems is via orthogonalization, transforming the agent’s future information so that it is independent of their current information. In this case, EsoSzentes2017 shows dynamic irrelevance in a different sense: For any implementable allocation rule, making the agent’s future orthogonalization information publicly observable does not change the monopolist’s profits. Does that result hold in the multidimensional setting and can it be operationalized?
To model the orthogonalization approach in the multidimensional setting, let $v(\gamma,z)$ be some function such that for $z\sim U([0,1]^{n})$ we have $v(\gamma,z)\sim F(\theta|\gamma)$ so $\theta=v(\gamma,z)$. This is similar a copula but dependencies between dimensions are captured in $v$ itself instead of in a separate function $C$; the model and results all hold using a copula instead. One such mapping is
$v(\gamma,z)=\left((F^{1})^{-1}(z^{1}|\gamma),(F^{2})^{-1}(z^{2}|\theta^{1},\gamma),...,(F^{n})^{-1}(z^{n}|\theta^{1},...,\theta^{n-1},\gamma)\right).$
In this setting (call it the relaxed problem), the monopolist’s problem is to choose $\hat{q}:\Gamma\times[0,1]^{n}\to[0,1]^{n},\hat{t}:\Gamma\times[0,1]^{n}\to\mathbb{R}$. Since screening essentially only happens once, only one transfer rule $\hat{t}$ is needed, instead of $t_{1},t_{2}$ as before. As such, the monopolist’s problem is:
$\max_{\hat{t},\hat{q}}$ $\>\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]$
s.t. $\>\gamma\in\operatorname*{\arg\!\max}_{\gamma^{\prime}}\mathbb{E}_{z}\left[\hat{q}(\gamma^{\prime},z)\cdot v(\gamma,z)-\hat{t}(\gamma^{\prime})\right];$ $IC^{\prime}$
$\>\mathbb{E}_{z}\left[\hat{q}(\gamma,z)\cdot v(\gamma,z)\right]-\hat{t}(\gamma)\geq 0\text{ for all }\gamma\in\Gamma$ $IR^{\prime}$
Two notes about this formulation. First, it would also be possible to write down a model where $\hat{t}$ also depends on the realized $z$, but by the same pushforward intuition as earlier, we can take $\hat{t}(\gamma)=\mathbb{E}_{z}\left[\hat{t}(\gamma,z)\right].$ Second, it is inconsequential if $z$ were observed before or after contracting; $IC^{\prime}$ could be replaced with a pointwise ($z$-by-$z$) incentive constraint instead. This case could be interpreted as a static multidimensional screening problem, but the buyer’s private information can be projected down to a one-dimensional parameter.
By similar reasoning as Lemma 1, we are able to recover the irrelevance result of EsoSzentes2017, even when there is dependence between dimensions (recall that Lemma 1 did not require independence either). However, Lemma 1 clarifies the mechanism by which post-contractual information rents need not be paid: they have simply already been extracted in period one. On the other hand, the reasoning of EsoSzentes2017 directly jumps to the case where no information rents need to be paid.
Proposition 2 (No Post-Contractual Information Rents).
Suppose $q$ is implementable in the original problem. Then, the monopolist’s profits from $q$ in the original problem are equal to their profits in the relaxed problem.
However, like the original versus sequential case, Proposition 2 does not imply that the original and relaxed problems are the same since implementability is an issue; more mechanisms can be implemented in the relaxed problem. When the distribution over values satisfies invariant dependencies, this gap disappears and the solution to the relaxed problem coincides to the solution of the original problem.
Proposition 3 (Relaxed Solution is Implementable).
If $F$ has invariant dependencies, then the solution of the original problem coincides with the solution of the relaxed problem.
However, invariant dependencies is violated, solutions to the relaxed problem are not implementable in the original problem and cyclic monotonicity becomes the key binding constraint; in a sequential sequential problem, cyclic monotonicity does not have any bite as contracting is done one dimension at a time.
4 Discussion
Endowing a principle the ability to contract intertemporally enables frontloading surplus extraction. This distorts the information rents which need to be paid to the agent in favor of the principle, increasing profits and simplifying the multidimensional screening problem when distributions satisfy invariant dependencies. In this case, the distortion of information rents enables the monopolist’s problem to be converted into one that is separable across dimensions; this transformation solves a probability continuity equation and corresponding to statistical tests of the agent’s reported pre-contractual information being separable across dimensions. When this is the case, the optimal mechanism sells goods separately, in contrast to complicated non-linear bundling usually featured in multidimensional screening problems. However, this mechanism parallels membership schemes retailers often employ in practice. Finally, these results formalize new notions of dynamic irrelevance in dynamic contracting.
There are several avenues for further investigation. First, offering the optimal sequential screening mechanism independently for each good is still a feasible mechanism even when invariant dependencies does not hold: Are there more general conditions for which it is optimal? Away from the invariant dependencies case, the function $\mathbf{V}$ may no longer be the same. However, a more complicated $\mathbf{V}$ may still be independent across dimensions, leading to a different pointwise maximization solution; even if the pointwise maximum is not implementable due to violations of cyclic monotonicity, it may still be the case that contracting independently across goods remains optimal. Second is analyzing the welfare effects of endowing the principal the ability to screen through time. Compared to the case where the monopolist can only contract after valuations themselves are known to the consumer, the monopolist always makes more profits in the sequential case (ignoring the agent’s pre-contractual private information and offering the optimal static multidimensional screening mechanism is always feasible). On the other hand, consumer information rents are distorted downwards but since it costs the monopolist less to procure information rents, realized information rents (net of distortions) may be larger. Investigating both how total surplus changes and how the division of surplus changes would be of interest.
Appendix A: Omitted Proofs
PROOF OF LEMMA 1
Proof.
For any $u(\gamma,\theta)$, applying the envelope theorem to period two gives that the allocation rule must be $q(\gamma,\theta)=\nabla_{\theta}u(\gamma,\theta)$. As such, period two transfers are
$t_{2}(\gamma,\theta)=\theta\cdot\nabla_{\theta}u(\gamma,\theta)-u(\gamma,\theta).$
Let $U(\gamma^{\prime}|\gamma)$ be the interim utility a consumer with type $\gamma$ receives when reporting $\gamma^{\prime}$ and $U(\gamma)=U(\gamma|\gamma)$ be the interim utility of a type $\gamma$ consumer when reporting truthfully. Since $\gamma$ is not directly payoff-relevant and only affects payoffs by influencing the distribution of $\theta$, we have
$U(\gamma^{\prime}|\gamma)=\int_{\Theta}u(\gamma^{\prime},\theta)dF(\theta|\gamma)-t_{1}(\gamma^{\prime});U(\gamma)=\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)-t_{1}(\gamma).$
Then, by incentive compatibility $\gamma\in\operatorname*{\arg\!\max}_{\gamma^{\prime}}U(\gamma^{\prime}|\gamma)$ so the Envelope Theorem applies and local incentive constraints gives
$U_{\gamma}(\gamma)=U_{\gamma}(\gamma^{\prime}=\gamma|\gamma)=\int_{\Theta}u(\gamma,\theta)f_{\gamma}(\theta|\gamma)d\theta.$
By the Fundamental Theorem of Calculus, we have
$U(\gamma)=U(\underline{\gamma},\underline{\gamma})+\int_{\underline{\gamma}}^{\gamma}U_{\gamma}(\gamma^{\prime})d\gamma^{\prime}=U(\underline{\gamma},\underline{\gamma})+\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}.$
As usual, we can set $U(\underline{\gamma},\underline{\gamma})=0$. Equating this with the direct computation of $U(\gamma)$ and solving for transfers gives
$t_{1}(\gamma)=\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)-\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}.$
Thus, $t_{1},t_{2},q$ are all pinned down by $u$ so the designer only needs to maximize over $u$; the monopolist’s objective function can be written as
$\mathbb{E}_{\gamma}\left[t_{1}(\gamma)\right]+\mathbb{E}_{\gamma,\theta}\left[t_{2}(\gamma,\theta)\right]=$ $\int_{\Gamma}\left[\int_{\Theta}u(\gamma,\theta)dF(\theta|\gamma)-\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}\right]dG(\gamma)$
$+\int_{\Gamma}\int_{\Theta}\left[\theta\cdot\nabla_{\theta}u(\gamma,\theta)-u(\gamma,\theta)\right]dF(\theta|\gamma)dG(\gamma).$
Observe that the first term of $\mathbb{E}_{\gamma}\left[t_{1}(\gamma)\right]$ and the second term of $\mathbb{E}_{\gamma,\theta}\left[t_{2}(\gamma,\theta)\right]$ cancel out: Any information rents given to the consumer for their post-contractual information can be extracted, in expectation, before that information arrives. The remainder of the computation is standard, interchanging the two integrals with respect to $\gamma$ to recover the standard hazard rate of $G$ terms: Expected transfers are
$\int_{\Gamma}\left[-\int_{\underline{\gamma}}^{\gamma}\int_{\Theta}u(\gamma^{\prime},\theta)f_{\gamma}(\theta|\gamma^{\prime})d\theta d\gamma^{\prime}+\int_{\Theta}\theta\cdot\nabla_{\theta}u(\gamma,\theta)f(\theta|\gamma)d\theta\right]g(\gamma)d\gamma$
$=$ $\int_{\Gamma}\int_{\Theta}\Big[-u(\gamma,\theta)f_{\gamma}(\theta|\gamma)(1-G(\gamma))+\theta\cdot\nabla_{\theta}u(\gamma,\theta)f(\theta|\gamma)g(\gamma)\Big]d\theta d\gamma$
$=$ $\mathbb{E}_{\gamma,\theta}\left[\theta\cdot\nabla_{\theta}u(\gamma,\theta)-u(\gamma,\theta)\frac{f_{\gamma}(\theta|\gamma)}{f(\theta|\gamma)}\frac{1-G(\gamma)}{g(\gamma)}\right]$
as desired. ∎
PROOF OF THEOREM 1
Proof.
Suppose $F$ has invariant dependencies. Then, suppressing function arguments and using the fact that for any function $h(x),\frac{d\ln(h(x))}{dx}=h^{\prime}(x)/h(x)$, at every $\theta$ and $\gamma$ we have
$\nabla_{\theta}(\mathbf{V}f)$ $=\sum_{j}\left[f\frac{\partial\mathbf{V}^{j}}{\partial\theta^{j}}+\mathbf{V}^{j}\frac{\partial f}{\partial\theta^{j}}\right]=\sum_{j}\left[f\frac{\partial}{\partial\theta^{j}}\left(\frac{F_{\gamma}^{j}}{f^{j}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}f\frac{\partial\ln(f)}{\partial\theta^{j}}\right]$
$=\sum_{j}\left[f\left(\frac{f^{j}f_{\gamma}^{j}-F_{\gamma}^{j}f^{j}_{\theta_{j}}}{(f^{j})^{2}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}f\frac{\partial}{\partial\theta^{j}}\left(\ln(c(F^{1},...,F^{n}))+\sum_{\ell}\ln(f^{\ell})\right)\right]$
$=f\sum_{j}\left[\left(\frac{f^{j}f_{\gamma}^{j}-F_{\gamma}^{j}f^{j}_{\theta_{j}}}{(f^{j})^{2}}\right)+\frac{F_{\gamma}^{j}}{f^{j}}\frac{\partial\ln(c)}{\partial u^{j}}f^{j}+F_{\gamma}^{j}\frac{f^{j}_{\theta^{j}}}{(f^{j})^{2}}\right]$
$=f\sum_{j}\left[\frac{f_{\gamma}^{j}}{f^{j}}+F_{\gamma}^{j}\frac{\partial\ln(c)}{\partial u^{j}}\right]=f\frac{\partial}{\partial\gamma}\left[\sum_{j}\ln(f^{j})+\ln(c(F^{1},...,F^{n}))\right]$
$=f\frac{\partial\ln(f)}{\partial\gamma}=f\frac{f_{\gamma}}{f}=f_{\gamma}.$
Then for every $\gamma,\theta$, the product rule gives
$\nabla_{\theta}\cdot\left(u(\gamma,\theta)\mathbf{V}(\theta|\gamma)f(\theta|\gamma)\right)$ $=\nabla_{\theta}u(\gamma,\theta)\cdot[\mathbf{V}(\theta|\gamma)f(\theta|\gamma)]+u(\gamma,\theta)[\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))]$
so plugging the above into the following integral and applying the Divergence Theorem gives
$\int_{\Theta}-u(\gamma,\theta)f_{\gamma}(\theta|\gamma)d\theta$ $=\int_{\Theta}-u(\gamma,\theta)[\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))]d\theta$
$=\int_{\Theta}\nabla_{\theta}u(\gamma,\theta)\cdot[\mathbf{V}(\theta|\gamma)f(\theta|\gamma)]-\nabla_{\theta}\cdot\left(u(\gamma,\theta)\mathbf{V}(\theta|\gamma)f(\theta|\gamma)\right)d\theta$
$=\int_{\Theta}\nabla_{\theta}u(\gamma,\theta)\cdot[\mathbf{V}(\theta|\gamma)f(\theta|\gamma)]d\theta-\oint_{\partial\Theta}(u(\gamma,\theta)\mathbf{V}(\theta|\gamma)f(\theta|\gamma))\cdot\mathbf{n}\>dS.$
Then, along the boundary $\partial\Theta$, for any $j$ $F^{j}(\theta^{j}|\gamma)$ is either zero or one for all $\gamma$. As such, $F^{j}_{\gamma}(\theta^{j}|\gamma)=0$ along the boundary and thus
$(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))^{j}=F^{j}_{\gamma}(\theta^{j}|\gamma)\prod_{i\neq j}f^{i}(\theta^{i}|\gamma)=0$
so the boundary term vanishes for any $u$. Thus, we are left with
$\int_{\Theta}-u(\gamma,\theta)f_{\gamma}(\theta|\gamma)d\theta=\int_{\Theta}\nabla_{\theta}u(\gamma,\theta)\cdot[\mathbf{V}(\theta|\gamma)f(\theta|\gamma)]d\theta;$
substituting this into the objective of Lemma 1 and using $\nabla_{\theta}u(\gamma,\theta)=q(\gamma,\theta)$ gives that the principal seeks to maximize
$\mathbb{E}_{\gamma,\theta}\left[\sum_{j}q^{j}(\gamma,\theta)\phi^{j}(\gamma,\theta^{j})\right].$
Maximizing pointwise gives the optimal allocation rule to be
$q(\gamma,\theta)=\sum_{j=1}^{n}e^{j}I\left\{\phi^{j}(\gamma,\theta^{j})\geq 0\right\}$
where $e^{j}$ is the $j$th basis vector and $I(A)$ is an indicator function over the set $A$. As virtual values are single-crossing in $\theta^{j}$ for each $j$ there is some threshold $\hat{\theta}^{j}(\gamma)$ such that
$\phi^{j}(\gamma,\theta^{j})\geq 0\iff\theta^{j}\geq\hat{\theta}^{j}(\gamma).$
Thus, $q(\gamma,\theta)$ is cyclically monotone in $\theta$ for each $\gamma$ and thus induces a utility function $u(\gamma,\theta)$ that is 1-Lipschitz and convex in $\theta$. As virtual values are increasing in $\gamma$, the threshold $\hat{\theta^{j}}(\gamma)$ is decreasing in $\gamma$ so $q^{j}(\gamma,\theta)$ is increasing in $\gamma$. Thus, first-period incentive constraints hold as well. ∎
PROOF OF PROPOSITION 1
Proof.
Let $t(\gamma,\theta)$ be net transfers. First, any $q(\gamma,\theta),t(\gamma,\theta)$ that is implementable in the original setting is implementable in the sequential setting. We will prove the contrapositive: If $q(\gamma,\theta),t(\gamma,\theta)$ is not implementable in the sequential setting, then it must not be implementable in the original setting.
Suppose $q(\gamma,\theta),t(\gamma,\theta)$ is not implementable in the sequential setting. Then, the buyer must have a profitable deviation at some point along the path of play of the mechanism. For notational simplicity, suppose the deviation happens after $\gamma$ has been truthfully reported; the same reasoning holds for deviations in $\gamma$ as well. Thus, there exists some $\gamma,\theta^{1},...,\theta^{k}$ such that the buyer has learned $\gamma,\theta^{1},...,\theta^{k}$, truthfully reported $\gamma,\theta^{1},...,\theta^{k-1}$, and can profitably deviate to a sequence of misreports $\{\tilde{\theta}^{j}(\theta_{1},...,\theta_{j})\}_{j=k}^{n}$. Slightly abusing notation, let $\tilde{\theta}(\theta)$ be the agent’s reports for a given realization of $\theta$. For this to be profitable, we have
$\mathbb{E}_{\theta^{k+1},...,\theta^{n}|\gamma}\left[\theta\cdot q(\gamma,\tilde{\theta}(\theta))-t(\gamma,\tilde{\theta}(\theta))\bigg|\theta^{1},..,\theta^{k}\right]>\mathbb{E}_{\theta^{k+1},...,\theta^{n}|\gamma}\left[\theta\cdot q(\gamma,\theta)-t(\gamma,\theta)\bigg|\theta^{1},..,\theta^{k}\right].$
A necessary condition for the above inequality to hold is for there to exist a realization $\theta$ for which
$\theta\cdot q(\gamma,\tilde{\theta}(\theta))-t(\gamma,\tilde{\theta}(\theta))>\theta\cdot q(\gamma,\theta)-t(\gamma,\theta).$
However, this implies that a buyer with realized values $\gamma,\theta$ as above has a profitable deviation to report $\gamma,\tilde{\theta}(\theta)$, as desired.
Then, the monopolist’s objective in the sequential case, as formalized in segal2014, coincides with the monopolist’s objective in the original setting, albeit with different constraints. Since pointwise maximization is implementable in the original problem, it is also implementable in the sequential case. ∎
PROOF OF PROPOSITION 2
Proof.
Suppose $q,t_{1},t_{2}$ is incentive compatible in the original problem and let $\hat{q}(\gamma,z)=q(\gamma,v(\gamma,z)),\hat{t}(\gamma)=\mathbb{E}_{z}\left[t_{1}(\gamma)+t_{2}(\gamma,v(\gamma,z))\right]$. Let
$\hat{U}(\gamma^{\prime}|\gamma)=\mathbb{E}_{z}\left[\hat{q}(\gamma^{\prime},z)\cdot v(\gamma,z)\right]-\hat{t}(\gamma^{\prime}),\hat{U}(\gamma)=\hat{U}(\gamma|\gamma)$
be the buyer’s interim utility function from reporting $\gamma^{\prime}$ at $\gamma$ and their interim utility when truthful. By the Envelope Theorem and interchanging differentiation and expectation,
$\hat{U}_{\gamma}(\gamma)=\mathbb{E}_{z}\left[\hat{q}(\gamma,z)\cdot v_{\gamma}(\gamma,z)\right].$
Then, applying the Fundamental Theorem of Calculus gives
$\mathbb{E}_{z}\left[\hat{q}(\gamma,z)\cdot v(\gamma,z)\right]-\hat{t}(\gamma)=\hat{U}(\gamma)=\hat{U}(\underline{\gamma})+\int_{\underline{\gamma}}\hat{U}_{\gamma}(\gamma^{\prime})d\gamma^{\prime}$
so transfers are now pinned down by (setting $\hat{U}(\underline{\gamma})=0$):
$\hat{t}(\gamma)$ $=\mathbb{E}_{z}\left[\hat{q}(\gamma,z)\cdot v(\gamma,z)-\int_{\underline{\gamma}}^{\gamma}\hat{q}(\gamma^{\prime},z)\cdot v_{\gamma}(\gamma^{\prime},z)d\gamma^{\prime}\right].$
Using the standard technique of interchanging the two integrals taken with respect to $\gamma$, expected transfers can be written as
$\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]=\mathbb{E}_{\gamma,z}\left[\hat{q}(\gamma,z)\cdot\left(v(\gamma,z)+v_{\gamma}(\gamma,z)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$
Then, $z=v^{-1}(\gamma,\theta)$ so taking $q(\gamma,\theta)=\hat{q}(\gamma,v^{-1}(\gamma,\theta))$ and $\mathbf{V}(\theta|\gamma)=v_{\gamma}(\gamma,v^{-1}(\gamma,\theta))$ gives expected transfers to be
$\mathbb{E}_{\gamma}\left[\hat{t}(\gamma)\right]=\mathbb{E}_{\gamma,\theta}\left[q(\gamma,\theta)\cdot\left(\theta+\mathbf{V}(\theta|\gamma)\frac{1-G(\gamma)}{g(\gamma)}\right)\right].$
As $q(\gamma,\theta)$ were implementable in the original problem, $q(\gamma,\theta)=\nabla_{\theta}u(\gamma,\theta).$ Then, expected profits as formulated above coincides with Equation 2 so it suffice to show that $\mathbf{V}(\theta|\gamma)$ satisfies $\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma))=f_{\gamma}(\theta|\gamma)$ and boundary conditions vanish.
By construction, $v$ satisfies
$F(v(\gamma,z)|\gamma)=\prod_{j}z^{j}$
so differentiating both sides with respect to $\gamma$ and replacing $z$ with $\theta$ gives
$\nabla_{\theta}F(v(\gamma,z)|\gamma)\cdot v_{\gamma}(\gamma,\theta)+F_{\gamma}(v(\gamma,z)|\gamma)=0\implies F_{\gamma}(\theta|\gamma)=-\sum_{i=1}^{n}\frac{\partial F}{\partial\theta^{i}}\mathbf{V}^{i}(\gamma,\theta).$
Using this identity, we have
$f_{\gamma}(\theta|\gamma)$ $=\frac{\partial^{n}}{\partial\theta^{1},...,\partial\theta^{n}}\left[F_{\gamma}(\theta|\gamma)\right]=\frac{\partial^{n}}{\partial\theta^{1},...,\partial\theta^{n}}\left[-\sum_{i=1}^{n}\frac{\partial F}{\partial\theta^{i}}\mathbf{V}^{i}(\gamma,\theta)\right]=-\nabla_{\theta}\cdot(\mathbf{V}(\theta|\gamma)f(\theta|\gamma)).$
Finally, boundary conditions vanish as for $\theta\in\partial\Theta$, we have that $v^{-1}(\gamma,\theta)\in\partial[0,1]^{n}$ and if each entry of $z$ is either zero or one then
$v^{j}(\gamma,z)=\begin{cases}\underline{\theta}^{j}&\text{ if }z^{j}=0;\\ \bar{\theta}^{j}&\text{ if }z^{j}=1.\end{cases}$
As such, $v(\gamma,z)$ is constant in $\gamma$ so $v_{\gamma}(\gamma,z)=0$ on $\partial[0,1]^{n}$ and boundary terms vanish. ∎
PROOF OF PROPOSITION 3
Proof.
Suppose $F$ satisfies invariant dependencies. Then, we can write
$\theta=F^{-1}(C(F^{1}(\theta^{1}|\gamma),...,F^{n}(\theta^{n}|\gamma))|\gamma)=F^{-1}(\eta|\gamma)$
for $\eta\sim C$. We also have $\theta=v(\gamma,z)$. Let $\kappa$ be a copula so $\eta=\kappa(z)$. Then,
$v^{j}(\gamma,z)=(F^{j})^{-1}(\kappa^{j}(z)|\gamma)$
and
$v^{-1}(\gamma,\theta)=z=\kappa^{-1}((F^{1})^{-1}(\theta^{1}|\gamma),...,(F^{n})^{-1}(\theta^{n}|\gamma))$
which also implies
$\kappa^{j}(v^{-1}(\gamma,\theta))=F^{j}(\theta^{j}|\gamma).$
Then,
$\mathbf{V}^{j}(\gamma,\theta)$ $=v_{\gamma}^{j}(\gamma,v^{-1}(\gamma,\theta))=\frac{\partial}{\partial\gamma}\left[v^{j}(\gamma,z)\right]_{z=v^{-1}(\gamma,\theta)}$
$=\frac{\partial}{\partial\gamma}\left[(F^{j})^{-1}(\kappa^{j}(z)|\gamma)\right]_{z=v^{-1}(\gamma,\theta)}$
$=\frac{\partial}{\partial\gamma}\left[(F^{j})^{-1}(\eta^{j}|\gamma)\right]_{\eta^{j}=\kappa^{j}(v^{-1}(\gamma,\theta))}$
$=\frac{\partial}{\partial\gamma}\left[(F^{j})^{-1}(\eta^{j}|\gamma)\right]_{\eta^{j}=F^{j}(\theta^{j}|\gamma)}$
$=-\frac{F^{j}_{\gamma}(\theta^{j}|\gamma)}{f^{j}(\theta^{j}|\gamma)}$
where the last line comes from the inverse function theorem. Thus, the objective found via Proposition 1 is equivalent to the objective found via Theorem 1 and the relaxed problem thus has the same solution as the original problem. Thus, the pointwise optimal allocation rule solves both problems. ∎